OpenStudy (anonymous):

f(θ) = (3cosθ−sinθ)^2

OpenStudy (anonymous):

simplify the espression

OpenStudy (mertsj):

\[(3\cos \theta-\sin \theta)^2=9\cos ^2\theta-6\sin \theta \cos \theta+\sin ^2\theta\]

OpenStudy (mertsj):

Remember that sin^2(x) + cos^2(x)=1

OpenStudy (anonymous):

1. f(θ) = 5−3sin2θ+4cos2θ 2. f(θ) = 4−3sin2θ+5cos2θ 3. f(θ) =−3+5sin2θ+4cos2θ 4. f(θ) = 4+4sin2θ−3cos2θ 5. f (θ) = 5+4sin2θ−3cos2θ 6. f(θ) = 3+4sin2θ+5cos2θ

OpenStudy (anonymous):

its one of those answers

OpenStudy (anonymous):

but idk which

OpenStudy (anonymous):

6sinθcosθ = 3sin2θ - thats one bit of it done

OpenStudy (mertsj):

1st one

OpenStudy (anonymous):

ahh yes 8 cos^2 θ + cos^2θ + sin^2θ = 8cos^2θ + 1 and 8 cos^2 θ = 4cos2θ + 4 combine to give 1st one

OpenStudy (mertsj):

\[9\cos^2\theta=9(1-\sin^2\theta)=9-9\sin^2\theta\] Now add the sin^2 theta \[9-9\sin^2\theta+\sin^2\theta=9+8\sin^2\theta\]

OpenStudy (mertsj):

But \[\sin^2\theta=\frac{1-\cos2\theta}{2}\]

OpenStudy (mertsj):

Whoops. Sign error 9-8sin^2 theta

OpenStudy (anonymous):

yup - very nice